OpenStudy (anonymous):

I have a triangular prism, and I have all the measurments of the base shape, except the length of the entire triangular prism. I also have the volume. Is there a certain formula, or do you have to rearrange one?

OpenStudy (lgbasallote):

yup there is

OpenStudy (lgbasallote):

\[V = \frac 13 \times (\frac{bh}{2} ) \times l\] where v = volume b = base of triangle h is height of triangle l is the length of the prism

OpenStudy (lgbasallote):

lol take away that 1/3

OpenStudy (anonymous):

Yea but I dont have the length, so what would the rearranged formula?

OpenStudy (anonymous):

 The side with the x, I dont have, but the side with the ticks I do

OpenStudy (lgbasallote):

you have to manipulate the formula i gave

OpenStudy (lgbasallote):

\[V = \frac{bh}{2} \times l\] \[2V = bhl\] \[\frac{2V}{bh} = l\]

OpenStudy (anonymous):

Thanks alot

OpenStudy (lgbasallote):

you're welcome